jacobrios15:

Which of the following shows the factors of 5x2 + 17x + 6? a-(5x + 1)(x + 6) b-(5x + 2)(x + 3) c-(5x + 3)(x + 2) d-(5x + 6)(x + 1)

Angle:

Hello! Welcome :) Give me a minute to read your question ^_^

jacobrios15:

ok:)

Angle:

hmm have you learned the FOIL method before?

jacobrios15:

no

Angle:

That's ok :) so the idea we can try here is called the Distributive property. a simple version of this idea is like this: 2*7 = 14 we know is true 2*(5+2) = 2*7 = 14 we know also works but we can also think of it as 2*(5+2) = 2*5 + 2*2 = 10 + 4 = 14 which also gets us to 14 so when working with variables, we can use this idea to make dealing with variables easier!

Angle:

Using this idea, we can look at Option A (5x + 1)(x+6) it can become = (5x + 1)*x + (5x + 1)*6 just like how the (2) in the example I gave, gets to be multiplied to both the (5) and the other (2)

Angle:

then just like how 2*5 = 10 = 5*2 the (5x + 1)*x we see above can be written as x*(5x + 1) then we can apply the Distributive Property again! x*(5x) + x*(1) then we simplify to get 5x^2 + x we can do the same to the other part: (5x + 1)*6 = 5x*(6) + 1*(6) = 30x + 6 then add both parts together (5x^2 + x) + (30x + 6) = 5x^2 + 31x + 6 This means that Option A: (5x + 1)(x+6) = 5x^2 + 31x + 6 but that is Not equal to 5x^2 + 17x + 6 so A is Not the answer we're looking for

Angle:

we can do a similar process with each option to get to your final answer ^_^

jacobrios15:

can you just give me the answer

Angle:

Personally, I would say that this Distributive Property thing is important and something that you will definitely have to use in a future math class. a-(5x + 1)(x + 6) = 5x^2 + 37x + 6 b-(5x + 2)(x + 3) = 5x^2 + 17x + 6 c-(5x + 3)(x + 2) = 5x^2 + 13x + 6 d-(5x + 6)(x + 1) = 5x^2 + 11x + 6

jacobrios15:

so whats the answer a,b,c,or d

Angle:

I have literally done all the work for you, compare the options with the given question 5x^2 + 17x + 6

jacobrios15:

is it b

Angle:

yup :)

jacobrios15:

gracias:)